# Functional Liftings of Restricted Geometric Inequalities

### v1

@@ 6. An Abstract Approach to the Equivalence of Geometric and Functional Inequalities
Numerous Brunn-Minkowski inequalities have been suggested in the setting of groups. In particular, for the Heisenberg group in [56, 39, 5], and extentions thereof [9, 67, 58]. Our machinery shows Tao’s Prékopa-Leindler inequality [67, Theorem 4] for nilpotent groups is equivalent to its set theoretic formulation, and consequently to the following. Given a simply connected nilpotent Lie group $G$ and subsets $A,B$, we define their product set to be
$A\cdot B=\{x\cdot y\colon x\in A,y\in B\},$
where $\cdot$ denotes the group product operation. For more details nilpotent Lie groups see [67, 58].
Theorem 6.10.
$h(x\cdot y)\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
holds for all $x,y$, then
where $\alpha^{\prime}=\frac{\alpha}{1+\alpha d}$.
Proof.
By [67, Theorem 3] or [58, Theorem 1.1], we see that
(22) $\mu^{\frac{1}{d}}(A\cdot B)\geq\mu^{\frac{1}{d}}(A)+\mu^{\frac{1}{d}}(B).$
Taking
$h(x\cdot y)$ $\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
an analogous but weaker inequality in the case that $\alpha\leq 1$.
Acknowledgements
This project was inspired in part by Gautam Aishwarya’s enthusiasm for Theorem 1.1 and its proof. The second author thanks Aishwarya and HIM in Bonn Germany for hosting these discussions during the Dual Trimester Program “Boolean Analysis in Computer Science”. He thanks Tomasz Tkocz for encouragement and helpful discussions as well.
Malliaris and Roberto are supported by the FP2M federation (CNRS FR 2036); this work received support from the Graduate School EUR-MINT (State support managed by the National Research Agency for Future Investments program bearing the reference ANR-18-EURE-0023).
Melbourne is supported by CONAHCYT Grant CBF2023-2024-3907.
Roysdon is supported by US NSF Grant DMS-2452384.

### v3

@@ 5. An Abstract Approach to the Equivalence of Geometric and Functional Inequalities
Numerous Brunn-Minkowski inequalities have been suggested in the setting of groups. In particular, for the Heisenberg group in [48, 32, 5], and extensions thereof [6, 56, 49]. Our machinery shows Tao’s Prékopa-Leindler inequality [56, Theorem 4] for nilpotent groups is equivalent to its set theoretic formulation, and consequently to the following. Given a simply connected nilpotent Lie group $G$ and subsets $A,B$, we define their product set to be
$A\cdot B=\{x\cdot y\colon x\in A,y\in B\},$
where $\cdot$ denotes the group product operation. For more details on nilpotent Lie groups see [56, 49].
Theorem 5.9.
$h(x\cdot y)\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
holds for all $x,y$, then
where $\alpha^{\prime}=\frac{\alpha}{1+\alpha d}$.
Proof.
By [56, Theorem 3] or [49, Theorem 1.1], we see that
(26) $\mu^{\frac{1}{d}}(A\cdot B)\geq\mu^{\frac{1}{d}}(A)+\mu^{\frac{1}{d}}(B).$
Taking
$h(x\cdot y)$ $\geq\mathcal{M}_{\alpha}^{(t)}(f(x),g(y))$
an analogous but weaker inequality in the case that $\alpha\leq 1$.
@@
Acknowledgments
This project was inspired in part by Gautam Aishwarya’s enthusiasm for Corollary 1.2 and its proof. The second author thanks Aishwarya and HIM in Bonn Germany for hosting these discussions during the Dual Trimester Program “Boolean Analysis in Computer Science”. He thanks Tomasz Tkocz for encouragement and helpful discussions as well. We would like to thank Franck Barthe for pointing out to us Uhrin’s reference [57], and Alexandros Eskenazis for bringing to our attention an error in an early version of this note and for helpful comments; in particular, for bringing to our attention how our results are connected to [16, Question 19].
Malliaris and Roberto are supported by the FP2M federation (CNRS FR 2036); this work received support from the Graduate School EUR-MINT (State support managed by the National Research Agency for Future Investments program bearing the reference ANR-18-EURE-0023). Melbourne is supported by CONAHCYT Grant CBF2023-2024-3907. Roysdon is supported by US NSF Grant DMS-2548742 (formerly classified as US NSF Grant DMS-2452384).